OpenStudy (anonymous):

Which of the binomials below is a factor of this trinominal? X^2-14x+24 X+6 X+4 X-12 X^2+6

OpenStudy (anonymous):

What numbers when multiplied together give you 24, but when added together would give you 14x?

OpenStudy (anonymous):

10

OpenStudy (the_fizicx99):

Correction, when you multiply them together, they give you 24, but when added together would give you -14

OpenStudy (anonymous):

Correct it is -14. Thanks

OpenStudy (the_fizicx99):

Make a list: 1 * 24 = 24 | 1 + 24 = 25 -1 * -24 = 24 | -1 - 24 = -25 2 * 12 = 24 | 2 + 12 = 14 -2 * -12 = 24 | -2 - 12 = -14 <-- And so on... So you have x^2 - 12x - 2x + 24, now factor this polynomial by grouping. :D

OpenStudy (anonymous):

So ot's c

OpenStudy (the_fizicx99):

Well there's an easy trick, you have x^2 So you'll have: (x )(x ) Plug in the terms that you solved, here -12 and -2 (x - 12)(x -2) Check: x^2 -2x - 12x + 24 = x^2 - 14x + 24 :D so this does work :> So yes, it's C :P

OpenStudy (anonymous):

Thank you :)

OpenStudy (the_fizicx99):

You're very welcome :P